The oxidation of ammonia by oxygen to form nitrogen and water releases a considerable amount of energy.
1.Identify the reactants.
2.Identify the products.
3.Write the balanced chemical equation.
Use the lowest whole-number coefficients possible, and include states of matter
4. 141.0 g of NH3 is mixed with 234.1 g of O2 and allowed to react. The reaction is complete when the NH3 is consumed. 116.0 g of N2 is formed and 35.40 g O2 remain unreacted. What mass of H2O is produced?
1
Expert's answer
2019-11-19T07:59:57-0500

Reactant NH3 and O2

Products are N2 and H2O

Balanced Equation is

4NH3 + 3O2= 2N2 + 6H2O

From the balanced reaction it can be said that when 96 gm of O2 reacted 108 gm of H2O is formed.

Here (234.1-35.40)gm= 198.70 gm of O2 reacted

Therefore the amount of H2O form = (108/96)*198.70

= 223.5gm